OpenStudy (anonymous):

What is the radius of a cylinder with a volume of 432 pi cm^3 if the height of the cylinder is equal to its diameter?

OpenStudy (campbell_st):

the cylinder has a radius r and height 2r then using the volume for\[V=\pi r^2h\] \[432\pi = \pi r^2(2r)\] or \[432\pi = 2\pi r^3\] divide by 2 pi \[216 = r^3\] you should be able to solve it from here

OpenStudy (anonymous):

Ah! That's fantastic! Thank you soooo much!